OpenStudy (anonymous):

Solve the equation for x: 5x + 2y = 15.

OpenStudy (theviper):

any other \(\Large{\cal{Equation}}\) ???

OpenStudy (theviper):

no then 5x+2y=15 =>5x=15-2y =>x=\(\LARGE{\frac{15-2y}{5}}\)

OpenStudy (anonymous):

omg im really sorry went to get a glass of water <3

OpenStudy (anonymous):

but thank you very much